# Discovering Lie Groups with Flow Matching

### v1

@@ 5.3 Quantifying discovered symmetries
Table 1: Wasserstein-1 distance between generated group elements and ground truth group elements in 2D experiments (lower is better).
Method $\mathrm{SO}(2)\to C_{4}$ $\mathrm{GL}(2,\mathbb{R})^{+}\to C_{4}$ $\mathrm{GL}(2,\mathbb{C})\to D_{4}$
LieGAN (1 gen.) $1.688$ $1.688$ $1.803$
LieGAN (2 gen.) $1.565$ $1.565$ $1.701$
LieGAN (4 gen.) $1.672$ $1.672$ $1.131$
LieFlow(Ours) $\mathbf{0.072}$ $\mathbf{1.321}$ $\mathbf{1.093}$
@@
Table 2: Wasserstein-1 distance between generated group elements and ground truth group elements in 3D experiments (lower is better).
Method $\mathrm{SO}(3)\to\mathrm{Tet}$ $\mathrm{SO}(3)\to\mathrm{Oct}$ $\mathrm{SO}(3)\to\mathrm{SO}(2)$
LieGAN (1 gen.) $2.320$ $2.303$ $0.0975$
LieGAN (3 gen.) $2.326$ $2.198$ $0.0693$
LieGAN (6 gen.) $2.105$ $1.682$ $0.0528$
LieFlow(Ours) $\mathbf{0.0885}$ $\mathbf{0.0962}$ $\mathbf{0.0429}$
5.4 Analysis on Last-Minute Mode Convergence
@@
Given the challenges in learning the octahedral group, we hypothesize that using flow matching for symmetry discovery is in fact quite a difficult task. We illustrate with an even simpler scenario of flow matching directly on group elements (matrices), from $\mathrm{SO}(2)$ to $\mathrm{C_{4}}$. As the target vector field $u_{t}$ and the probability path $p_{t}(x_{t}\mid x_{1})$ are important quantities in our FM objective (equation 5), we visualize them for a single sample in Figure 6.
@@
To verify our hypothesis, we perform flow matching directly on the group elements from $\mathrm{SO}(2)$ to $C_{4}$. The group elements can be parameterized by a scalar $\theta$, allowing us to analyze the behavior of the learned vector field more easily. Each data sample $x_{1}$ is a scalar from the set $\{-\pi,-\pi/2,0,\pi/2\}$ and the source distribution is $q=\mathcal{U}[-\pi,\pi]$. We use the same network architecture as in the 2D experiments. We can further compute the reverse conditional or the posterior $p_{t}(x_{1}\mid x_{t})$ that can help verify two things: 1) that $x_{t}$ remain nearly stationary until $t$ approaches $1$, i.e. $p(x_{l}\mid x_{t})$ remains nearly uniform, and 2) $x_{t}$ chooses the closest group element $x_{1}$. The exact details of how $p_{t}(x_{1}\mid x_{t})$ is computed is given in Appendix D.
Figure 29 in Appendix C.6 visualizes the progression of $x_{t}$ over time for $100$ samples, showing that $C_{4}$ symmetry is clearly identified. Figure 7(a) plots the entropy of the posterior, and shows that even in this simple scenario, the entropy remains near the maximum (uniform distribution) until $t$ approaches $1$. Figure 7(b) visualizes the posterior $p(x_{1}\mid x_{t})$ averaged over 1,0001{,}000 samples and separated into $4$ classes, depending on the original $x_{0}$ and its nearest mode in $\{-\pi,-\pi/2,0,\pi/2\}$ (only two shown, other classes are similar, see Figure 30(b) in Appendix C.6). For all 4 classes, we observe that $p(x_{1}\mid x_{t})$ remains nearly uniform until around $t=0.9$, when it becomes uni-modal to the closest mode/component, e.g., for all $x_{0}$ that were closest to the mode $\mu=-\pi$, then the posterior $p(x_{1}\mid x_{t})$ converges to $\mu$. This shows both that the learned vector field remains nearly 0 for most timesteps, only converges when little time is left, and converges to the closest group element.
Note that this differs from the standard FM setting, where the flow is defined on a high-dimensional space. bertrand2025closed find that for high-dimensional flow matching with images, the entropy of the posterior drops at small times, transitioning from a stochastic phase that enables generalization to a non-stochastic phase that matches the target modes. In our low-dimensional setting of symmetry discovery, we find the opposite occurs: the entropy stays near uniform until large $t$, where it suddenly converges. We term this phenomenon “last-minute mode convergence”.
@@
Figure 8 shows the evolution of a uniform grid over $\mathrm{SO}(2)$ over time. As we hypothesized, we can see from the velocity arrows that particles move slowly closer to the midpoint between two modes ($t=0$ to $t=0.60$) but the velocities are relatively small. From $t=0.70$ onwards, we can see that the shift towards the closest modes and the velocities become increasingly large until $t=1$. Figure 31 (Appendix C.6) shows a clearer picture of how the velocity evolves over time. From $t=0$ to $t=0.6$, we see that the $x_{t}$’s are nudged towards the midpoints between modes. Interestingly, we see some oscillating behavior where the vector field flips sign, suggesting that the net velocity hovers close to 0 with no defined pattern. From $t=0.8$ onwards, we see that points $x_{t}$ are pushed to the closest $x_{1}$ mode, with increasing velocities. With near zero net velocity until $t=0.6$, these plots also support the observation of “last-minute mode convergence”.
6 Discussion
Despite the challenges revealed in our experiments, this work demonstrates that generative models, particularly flow matching, offer a promising avenue for symmetry discovery. We introduced a novel approach to symmetry discovery using flow matching over Lie groups, where we learn a map from a larger prior group to the data symmetries. Our experiments revealed both successes and fundamental challenges. In the 2D case, we successfully discovered discrete symmetries from several different continuous prior groups, including the dihedral group by learning a flow over the complex domain. The method naturally discovers the closest group element rather than exact transformations, which may be advantageous for handling approximate symmetries in real data. However, our 3D experiments exposed some limitations: while our method can discover smaller subgroups such as the tetrahedral group, it fails for higher order groups such as the octahedral group. Our analysis uncovered a ”last-minute convergence” phenomenon, that the symmetric arrangement of the target modes causes the learned vector field to remain near zero for most of the trajectory, only converging to the nearest mode as $t$ approaches $1$. Our simplified flow matching experiment supported this claim and we showed that the posterior entropy remains near uniform for most timesteps. We introduced a power time schedule to skew the sampling towards timesteps near $1$. Our heuristic allows use to discover the more complex octahedral group, but still fails for higher-order groups.
6.1 Limitations and Future Work
There are several limitations to our work. While we demonstrate feasibility on simple and noise-free point clouds, scaling to noisy real-world datasets remains an open question. Our method is currently limited to discovering symmetries within the parameterization of the hypothesis group, and it is still an open question as to whether LieFlow will scale to real-world datasets. Our method requires fitting a distribution over the prior group, which may impact the range of symmetries it can discover depending on the parameterization and boundaries. Although we have analyzed the cause behind last-minute mode convergence and proposed a time schedule according to the power distribution, it is still unclear how to discover complex symmetries. An important future direction is to investigate Riemannian diffusion–based formulations, which can sidestep the issue of discontinuities of the target vector field at the cut locus (huang2022riemannian; zhu2025trivialized; mangoubiefficient), and may alleviate the last-minute mode convergence behavior observed in LieFlow. Another important future direction is to improve numerical stability and scalability on high-dimensional or non-compact groups, enabling LieFlow to fully leverage its group-agnostic framework beyond the lower-dimensional settings demonstrated here.
Future directions include developing specialized architectures with group-aware inductive biases, incorporating additional supervision signals to guide convergence, and extending the method to handle approximate symmetries common in real-world data. The integration of discovered symmetries into downstream equivariant networks also remains an important open question. Despite current limitations, this work demonstrates that flow matching offers a promising path toward automatic symmetry discovery, lending us insights into the data and offering improved sample efficiency and generalization for downstream tasks.
Acknowledgements
The authors are grateful to Jianke Yang and Rose Yu for helpful discussions. F.E. and J.W.M. would like to acknowledge support from the Bosch Center for Artificial Intelligence. J.W.M. acknowledges additional support from the European Union’s Horizon Framework Programme (Grant agreement ID: 101120237). R.W. would like to acknowledge support from NSF Grants 2134178 and 2442658. L.L.S.W. would like to acknowledge support from NSF Grants 2107256.
LLM Usage
LLMs were used to revise sentences, correct grammar, and draft portions of the introduction, method, and discussion (Sections 1, 4 and 6). They were also used to generate parts of the model code and most of the visualization code.
Appendix A Derivation of the Target Flow Field
Interpolants on Lie Groups.
To derive FM on a Lie group, we need to define the interpolant as in the Euclidean case. An exponential curve starting at $g_{0},g_{1}\in G$ can be defined by (sherry2025flow):
$\gamma:[0,1]\to G;\quad t\mapsto g_{0}\exp(t\log(g_{0}^{-1}g_{1})).$ (6)
$u_{t}(g_{t}\mid g_{1})=\frac{(L_{g_{t}})_{*}\log(g_{t}^{-1}g_{1})}{1-t}.$ (7)
Defining the Target Vector Field.
Note that we couple source and target by an explicitly sampled group action. According to equation 4, given $x_{1}\sim q$ and $g\sim p(G)$, and set $x_{0}=gx_{1}$ and $A:=\log(g^{-1})\in\mathfrak{g}$. Consider the group curve $g_{t}=g\exp(tA)$ so that $g_{t}^{-1}=\exp((1-t)A)$ and the induced path $x_{t}=g_{t}x_{1}=x_{0}\exp(tA)$. Plugging $g_{1}=e$ into the group interpolant equation 7, we have:
$u_{t}(g_{t}\mid e)=\frac{(L_{g_{t}})*\log(g_{t}^{-1})}{1-t}=(L_{g_{t}})*A,$ (8)
Appendix B Training Details
(a) 2D arrow
(b) 2D half arrow
(c) 3D irregular tetrahedron
Figure 9: Canonical objects used for generating the 2D (a,b) and 3D (c) datasets
2D Datasets.
For the 2D point cloud datasets, we use canonical arrow and half-arrow objects as shown in Figure 9a,b. We generate datasets containing discrete symmetries, $C_{4}$ and $D_{4}$ transforming the canonical objects. The prior known groups are $\mathrm{SO}(2),\mathrm{GL}(2,\mathbb{R})^{+}$ for discovering $C_{4}$ and $\mathrm{GL}(2,\mathbb{C})$ for discovering $D_{4}$. For $\mathrm{SO}(2)$, we use a uniform distribution over the parameter $\theta\sim\mathcal{U}[-\pi,\pi]$ and generate the $2\times 2$ rotation matrices. For the $\mathrm{GL}(2)$ groups, we fit a pushforward distribution via the exponential map from a uniform distribution $\mathcal{U}[-\pi/2,\pi/2]$ over the Lie algebra coefficients and use the standard basis for the generators. On the network side, we use a small MLP and concatenate time $t$ to $x_{t}$ as inputs.
3D Datasets.
We use an irregular tetrahedron with no self-symmetries as shown in Figure 9c and generate datasets containing $\mathrm{Tet},\mathrm{Ico}$ symmetries. We consider the prior group $\mathrm{SO}(3)$ and use a uniform distribution over it by using Gaussian normalization over unit quaternions and transforming them into $3\times 3$ matrices.
For the 3D case, we use a similar but wider network architecture as the 2D datasets and use a sinusoidal time embedding.
Appendix C Results
This section contains additional results and visualizations for both 2D and 3D datasets.
C.1 2D Datasets
Figure 10: Histogram of $5000$ angles from the generated transformation matrices for $\mathrm{SO}(2)$ to $C_{4}$.
Figure 10 is the histogram of the angles of the generated transformation matrices for $\mathrm{SO}(2)$ to $C_{4}$. We can see that the model learns to generate angles close to multiples of $\pi/2$, corresponding to $C_{4}$ elements.

### v2

@@ 5.4 Results on Synthetic 3D Datasets
Table 2: Wasserstein-1 distance between generated group elements and ground truth group elements in irregular tetrahedron (Section 5.4) and rotated ModelNet10 (Section 5.5) experiments (lower is better). $\mathrm{SO}(3)\to\mathrm{SO}(2)$(a) corresponds to rotations around $z$ axis and $\mathrm{SO}(3)\to\mathrm{SO}(2)$(b) corresponds to rotations around $(0,1/2,-\sqrt{3}/2)$ axis. $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/o pow.) indicates results without power time schedule and $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/ pow.) indicates results with power time schedule.
(a) 3D irregular tetrahedron (Section 5.4)
Method $\mathrm{SO}(3)\to\mathrm{Tet}$ $\mathrm{SO}(3)\to\mathrm{Oct}$ $\mathrm{SO}(3)\to\mathrm{SO}(2)$(a) $\mathrm{SO}(3)\to\mathrm{SO}(2)$(b) $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/o pow.) $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/ pow.)
LieGAN $2.326$ $2.198$ $0.0693$ $0.0309$ $1.861$ $1.861$
LieFlow(Ours) $\mathbf{0.0885}$ $\mathbf{0.0962}$ $\mathbf{0.0429}$ $\mathbf{0.0299}$ $\mathbf{0.319}$ $\mathbf{0.118}$
(b) ModelNet10 (Section 5.5)
Method $\mathrm{SO}(3)\to\mathrm{Tet}$ $\mathrm{SO}(3)\to\mathrm{Oct}$ $\mathrm{SO}(3)\to\mathrm{SO}(2)$(a) $\mathrm{SO}(3)\to\mathrm{SO}(2)$(b) $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/o pow.) $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/ pow.)
LieGAN $2.291$ $2.233$ $\mathbf{0.0291}$ $0.0621$ $1.828$ $1.828$
LieFlow(Ours) $\mathbf{0.113}$ $\mathbf{0.125}$ $0.0461$ $\mathbf{0.0472}$ $\mathbf{0.717}$ $\mathbf{0.156}$
Since most of the learning signal for discrete groups occurs near $t=1$, we leverage a time scheduling scheme to sample more timesteps near $t=1$. Instead of sampling from $\mathcal{U}(0,1)$, we use a power distribution with density $p(x)=nx^{n-1},x\sim\mathcal{U}(0,1)$, where $n$ controls the skewness (see Figure 13 in Appendix E). We find empirically that $n=5$ works best; see Appendix E.1. Figure 14 in Appendix E confirms that with the power time schedule, objects begin moving earlier in the flow. With this power schedule, our model correctly learns the icosahedral group as shown in Figure 3(f).
5.5 Results on ModelNet10
Table 2(b) shows the quantitative results for different target symmetry groups of real-world ModelNet10 objects, with qualitative visualizations in Appendix F.3. As with the synthetic 3D datasets, LieFlow significantly outperforms LieGAN on all discrete groups and achieves comparable performance for continuous groups, demonstrating its effectiveness in discovering symmetries in real-world datasets. Again we find that using the power time schedule over uniform sampling improves performance for the higher-order discrete icosahedral group.
Incomplete Observations
To evaluate the robustness of LieFlow to incomplete observations, we randomly mask a portion of the points in ModelNet10 point clouds. Table 3 shows the results for the challenging setting of discovering $\mathrm{Ico}$ with hypothesis group $\mathrm{SO}(3)$ with different masking ratios (0%, 5%, 10%). Even with $10\%$ masking, LieFlow still discovers the underlying symmetries with only moderate degradation in performance. See Figures 22(g)22(h) in the Appendix F.3 for visualizations.
@@ 5.5 Results on ModelNet10
Table 3: Incomplete observations: Wasserstein-1 distance between generated group elements and ground truth group elements for $\mathrm{SO}(3)\to\mathrm{Ico}$ with different masking ratios (lower is better).
Method 0% 5% 10%
LieFlow 0.156 0.186 0.212
6 Conclusion and Future Work
Our work demonstrates that expressive generative models like flow matching offer a promising avenue for symmetry discovery. We introduced LieFlow, a novel framework that formulates symmetry discovery as learning a distribution on Lie groups using flow matching. By learning a flow from a large hypothesis group $G$ to a distribution whose support matches the true symmetry subgroup $H\subseteq G$, our approach discovers both continuous and discrete symmetries within a unified framework, without making assumptions on the Lie algebra basis or the distribution of its coefficients.
Our experiments demonstrate that LieFlow successfully discovers a variety of symmetry groups across different settings, such as $C_{4}$ and partial $C_{4}$ symmetries in 2D datasets, and several discrete and continuous $\mathrm{SO(3)}$ subgroups acting on 3D space. In addition, our method also demonstrates robustness to incomplete observations, with only a moderate decrease in performance.
Our study focuses on small point clouds and a limited number of objects, and we leave broader data modalities and larger-scale settings to future work. Moreover, our method requires fitting a distribution over the prior group, and the choice of the parameterization can impact performance, as seen with the two different prior distributions for $\mathrm{GL}(2)$. Important future directions include improving scalability to high-dimensional or non-compact groups, incorporating additional supervision signals to guide convergence, and integrating discovered symmetries into downstream equivariant networks. This work demonstrates that flow matching offers a promising path toward automatic symmetry discovery, lending us insights into the data and offering improved sample efficiency and generalization for downstream tasks.
Acknowledgements
The authors are grateful to Nima Dehmamy, Minghan Zhu and Rose Yu for helpful discussions. F.E. and J.W.M. would like to acknowledge support from the Bosch Center for Artificial Intelligence. J.W.M. acknowledges additional support from the European Union’s Horizon Framework Programme (Grant agreement ID: 101120237). L.L.S.W. would like to acknowledge support from NSF Grants 2107256. R.W. would like to acknowledge support from NSF Grants 2442658 and 2134178. This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
LLM Usage
LLMs were used to revise sentences, correct grammar, and draft portions of this work. They were also used to generate parts of the model code and most of the visualization code.
Appendix A Derivation of the Target Flow Field
Interpolants on Lie Groups.
To derive FM on a Lie group, we need to define the interpolant as in the Euclidean case. An exponential curve starting at $g_{0},g_{1}\in G$ can be defined by (Sherry and Smets, 2025):
$\gamma:[0,1]\to G;\quad t\mapsto g_{0}\exp(t\log(g_{0}^{-1}g_{1})).$ (7)
$u_{t}(g_{t}\mid g_{1})=\frac{(L_{g_{t}})_{*}\log(g_{t}^{-1}g_{1})}{1-t}.$ (8)
Defining the Target Vector Field.
Note that we couple source and target by an explicitly sampled group action. According to (4), given $x_{1}\sim q$ and $g\sim p(G)$, and set $x_{0}=gx_{1}$ and $A:=\log(g^{-1})\in\mathfrak{g}$. Consider the group curve $g_{t}=g\exp(tA)$ so that $g_{t}^{-1}=\exp((1-t)A)$ and the induced path $x_{t}=g_{t}x_{1}=x_{0}\exp(tA)$. Plugging $g_{1}=e$ into the group interpolant (8), we have:
$u_{t}(g_{t}\mid e)=\frac{(L_{g_{t}})*\log(g_{t}^{-1})}{1-t}=(L_{g_{t}})*A,$ (9)
Appendix B Dataset
(a) 2D arrow
(b) 3D irregular tetrahedron
Figure 5: Canonical objects used for generating the 2D (a) and 3D (b) datasets
Figure 6: Samples of the ModelNet10 rotated under $\mathrm{Tet}$ group.
2D arrow.
For the 2D synthetic datasets, we use canonical arrow object as shown in Figure 5a. We generate datasets with discrete symmetries $C_{4}$.
3D irregular tetrahedron.
We use an irregular tetrahedron with no self-symmetries as shown in Figure 5b and generate datasets containing $\mathrm{Tet},\mathrm{Oct},\mathrm{Ico}$ and $\mathrm{SO}(2)$ symmetries.
ModelNet10.
To construct the rotated ModelNet10 dataset, we randomly select 10 objects from each category of ModelNet10 (Wu et al., 2015) as canonical shapes and uniformly sample 128 points from each surface to form point clouds. Each data point is then generated by independently sampling a canonical shape and a transformation from the target group and applying the transformation to the point cloud. Figure 6 shows some samples of the ModelNet10 rotated under $\mathrm{Tet}$ group.
Appendix C Training Details
Sampling Groups from Prior Distribution.
For 2D arrow experiments, the prior known groups are $\mathrm{SO}(2),\mathrm{GL}(2,\mathbb{R})^{+}$. For $\mathrm{SO}(2)$, we use a uniform distribution over the parameter $\theta\sim\mathcal{U}[-\pi,\pi]$ and generate the $2\times 2$ rotation matrices. For the $\mathrm{GL}(2)^{+}$ groups, we use two different sample strategies:
  • Lie algebra coefficients sampling: We sample the coefficients $a,b,c,d\sim\mathcal{U}[-\pi,\pi]$ and construct the Lie algebra element as $A=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$. Then we exponentiate it to obtain the transformation matrix $g=\exp(A)$.
  • Matrix composition based sampling: We sample rotation angle $\theta\sim\mathcal{U}[-\pi,\pi]$, scaling factors $s_{x},s_{y}\sim\mathcal{U}[0.5,2.0]$ and shear factors $s_{h}\sim\mathcal{U}[-1,1]$. Then we construct the transformation matrix as $g=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}s_{x}&0\\ 0&s_{y}\end{bmatrix}\begin{bmatrix}1&s_{h}\\ 0&1\end{bmatrix}$.
Then, we reject any sampled transformation matrices with determinants outside $(0,2]$ to ensure they belong to $\mathrm{GL}(2)^{+}$ and make numerical stable.
For 3D experiments, we consider the prior group $\mathrm{SO}(3)$ and use a uniform distribution over it by using Gaussian normalization over unit quaternions and transforming them into $3\times 3$ matrices.
Training setting for Synthetic data
We use a time-conditioned MLP for flow matching that combines a sinusoidal time embedding module with a 5-layer MLP. The time embedding module uses sinusoidal positional encoding to transform the scalar time $t$ into a hidden dimensional vector, which is then projected to input dimension through a two-layer MLP with ReLU activations. The main network processes the concatenation of the flattened input and the time embedding through four hidden layers and an output layer with ReLU activations. The hidden dimension is set to $128$ for 2D arrow experiment and $512$ for 3D irregular tetrahedron experiment. We use the Adam optimizer with a learning rate of $3\times 10^{-3}$ for both experiments.
Training setting for ModelNet10
For the ModelNet10 experiment,the prediction network adopted is a two-layer Transformer encoder with four attention heads and a hidden dimension of $256$. To condition the network on the time $t$, we use $64$-dimensional sinusoidal time embeddings and concatenate them to the point features. The resulting per-point representations are aggregated via mean pooling to form a global feature. Then, the global feature is then passed through a three-layer MLP to predict a 3D Lie algebra element. AdamW optimizer is used with a learning rate of $1\times 10^{-4}$ and weight decay of $1\times 10^{-2}$.
Time comparison
We conduct training time comparisons between LieFlow and LieGAN on the Modelnet10 datasets. Both methods are trained on a single GeForce RTX 4090 GPU. Fixing batch size to $512$, for one epoch, LieFlow takes approximately $18$ seconds, while LieGAN takes around $50$ seconds.
Appendix D Analysis on Last-Minute Mode Convergence
@@
Given the challenges in learning the icosahedral group, we hypothesize that using flow matching for discovering discrete groups is in fact quite a difficult task. We illustrate with an even simpler scenario of flow matching directly on group elements (matrices), from $\mathrm{SO}(2)$ to $\mathrm{C_{4}}$. As the target vector field $u_{t}$ and the probability path $p_{t}(x_{t}\mid x_{1})$ are important quantities in our FM objective ((5)), we visualize them for a single sample in Figure 7.
@@
To verify our hypothesis, we perform flow matching directly on the group elements from $\mathrm{SO}(2)$ to $C_{4}$. The group elements can be parameterized by a scalar $\theta$, allowing us to analyze the behavior of the learned vector field more easily. Each data sample $x_{1}$ is a scalar from the set $\{-\pi,-\pi/2,0,\pi/2\}$ and the source distribution is $q=\mathcal{U}[-\pi,\pi]$. We use the same network architecture as in the 2D experiments. We can further compute the reverse conditional or the posterior $p_{t}(x_{1}\mid x_{t})$ that can help verify two things: 1) that $x_{t}$ remain nearly stationary until $t$ approaches $1$, i.e. $p(x_{l}\mid x_{t})$ remains nearly uniform, and 2) $x_{t}$ chooses the closest group element $x_{1}$. The exact details of how $p_{t}(x_{1}\mid x_{t})$ is computed is given in Appendix D.2.
Figures 10 in Appendix D.1 visualizes the progression of $x_{t}$ over time for $100$ samples, showing that $C_{4}$ symmetry is clearly identified. Figures 8(a) plots the entropy of the posterior, and shows that even in this simple scenario, the entropy remains near the maximum (uniform distribution) until $t$ approaches $1$. Figure 8(b) visualizes the posterior $p(x_{1}\mid x_{t})$ averaged over 1,0001{,}000 samples and separated into $4$ classes, depending on the original $x_{0}$ and its nearest mode in $\{-\pi,-\pi/2,0,\pi/2\}$ (only two shown, other classes are similar, see Figures 11(b) in Appendix D.1). For all 4 classes, we observe that $p(x_{1}\mid x_{t})$ remains nearly uniform until around $t=0.9$, when it becomes uni-modal to the closest mode/component, e.g., for all $x_{0}$ that were closest to the mode $\mu=-\pi$, then the posterior $p(x_{1}\mid x_{t})$ converges to $\mu$. This shows both that the learned vector field remains nearly 0 for most timesteps, only converges when little time is left, and converges to the closest group element.
Note that this differs from the standard FM setting, where the flow is defined on a high-dimensional space. Bertrand et al. (2025) find that for high-dimensional flow matching with images, the entropy of the posterior drops at small times, transitioning from a stochastic phase that enables generalization to a non-stochastic phase that matches the target modes. In our low-dimensional setting of symmetry discovery, we find the opposite occurs: the entropy stays near uniform until large $t$, where it suddenly converges. We term this phenomenon “last-minute mode convergence”.
@@
Figure 9 shows the evolution of a uniform grid over $\mathrm{SO}(2)$ over time. As we hypothesized, we can see from the velocity arrows that particles move slowly closer to the midpoint between two modes ($t=0$ to $t=0.60$) but the velocities are relatively small. From $t=0.70$ onwards, we can see that the shift towards the closest modes and the velocities become increasingly large until $t=1$. Figures 12 (Appendix D.1) shows a clearer picture of how the velocity evolves over time. From $t=0$ to $t=0.6$, we see that the $x_{t}$’s are nudged towards the midpoints between modes. Interestingly, we see some oscillating behavior where the vector field flips sign, suggesting that the net velocity hovers close to 0 with no defined pattern. From $t=0.8$ onwards, we see that points $x_{t}$ are pushed to the closest $x_{1}$ mode, with increasing velocities. With near zero net velocity until $t=0.6$, these plots also support the observation of “last-minute mode convergence”.
D.1 Flow Matching on Group Elements
This section contains additional visualizations for the flow matching experiment on group elements from $\mathrm{SO}(2)$ to $C_{4}$ as described in Section D.
Figures 10 shows the time progression of $x_{t}$ when generating $100$ samples from $t=0$ to $t=1$. We can see that the samples converge to the closest group element in the orbit.
Figures 11(b) in Figures 11 shows the posterior $p(x_{1}\mid x_{t})$ for all $4$ classes, depending on the original $x_{0}$ and its nearest mode in $\{-\pi,-\pi/2,0,\pi/2\}$. We can see that for all $4$ classes, the posterior $p(x_{1}\mid x_{t})$ remains nearly uniform when $t$ closes to $1$.
(a) $H_{p}(x_{1}\mid x_{t})$
(b) $p(x_{1}\mid x_{t})$
Figures 11 visualizes the posterior $p(x_{1}\mid x_{t})$, divided into 4 buckets depending on the initial position $x_{0}$. We see that the posterior remains nearly uniform until $t=0.95$, and then $x_{t}$ converges to the mode that was the closest when $t=0$.
Figures 12 visualizes how the velocities change over time. Until $t=0.6$, the velocities are close to 0 but slightly push the $x_{t}$’s towards the midpoints between modes. From $t=0.8$ onwards, we see that points $x_{t}$ are pushed to the closest $x_{1}$ mode, with increasing velocities.
D.2 Derivation of Posterior Computation
To compute the posterior, we use Bayes’ rule as
$p(x_{1}|x_{t})=\frac{p(x_{t}|x_{1})p(x_{1})}{p_{t}(x_{t})},$ (11)
We first sample initial points $x_{0}$ from a uniform prior over $\pi,\pi)$ and forward simulate using the learned velocity field $v_{\theta}$ to obtain the trajectories $x_{t}$ and their log probabilities through the continuity equation. To evaluate the likelihood $p(x_{t}\mid x_{1})$ for each candidate $x_{1}$, we invert the linear interpolation $x_{t}=(1-t)x_{0}+tx_{1}$ to find $x_{0}=(x_{t}-tx_{1})/(1-t)$, then integrate the divergence along the trajectory from this inverted $x_{0}$. Special handling is required at the boundaries: at $t=0$, the posterior equals the prior $p(x_{1})$ (which we know in this simple scenario) due to independence, while near $t=1$, we use a Gaussian approximation to avoid numerical instability from division by (1t)(1-t).
Algorithm 4 Compute Posterior $p(x_{1}\mid x_{t})$ for Flow Matching
0: Velocity field $v_{\theta}(x,t)$, prior $p_{0}\sim\mathcal{U}[-\pi,\pi)$, target $p_{1}(x_{1})=\sum_{k=1}^{K}\frac{1}{K}\delta(x_{1}-\mu^{(k)})$, time steps $T$
1:Initialize: Sample $\{x_{0}\}^{N}\sim p_{0}$
2:
3:Forward Simulation:
4:for $t\in\{0,\frac{1}{T},\ldots,1\}$ do
5:  $x_{t}\leftarrow\text{ODESolve}(v_{\theta},x_{0},[0,t])$ $\triangleright$ Integrate velocity field
6:  $\log p_{t}(x_{t})\leftarrow\log p_{0}(x_{0})-\int_{0}^{t}\nabla\cdot v_{\theta}(x_{s},s)ds$ $\triangleright$ Continuity equation
7:end for
8:
9:Posterior Computation:
10:for $t\in\{0,\frac{1}{T},\ldots,1\}$ do
11:  if $t=0$ then
12:   $p(x_{1}\mid x_{0})\leftarrow p_{1}(x_{1})$ $\triangleright$ Independent at $t=0$
13:  else if $t\approx 1$ then
14:   for $k=1$ to $K$ do
15:    $\log p(x_{t}\mid x_{1}^{(k)})\leftarrow-\frac{\|x_{t}-x_{1}^{(k)}\|^{2}}{2\sigma^{2}}$, where $\sigma=0.1$ $\triangleright$ Gaussian Approximation
16:   end for
17:  else
18:   for $k=1$ to $K$ do
19:    $x_{0}^{(k)}\leftarrow\frac{x_{t}-t\cdot x_{1}^{(k)}}{1-t}$ $\triangleright$ Invert linear interpolation
20:    Forward simulate path from $x_{0}^{(k)}$ to time $t$ using $v_{\theta}$ and continuity equation
21:    $\log p(x_{t}\mid x_{1}^{(k)})\leftarrow\log p_{0}(x_{0}^{(k)})-\int_{0}^{t}\nabla\cdot v_{\theta}(x_{s},s)ds$
22:   end for
23:  end if
24:  for $k=1$ to $K$ do
25:   $\log p(x_{1}^{(k)}\mid x_{t})\leftarrow\log p(x_{t}\mid x_{1}^{(k)})+\log p_{1}(x_{1}^{(k)})$ $\triangleright$ Bayes’ Rule
26:  end for
27:
28:  $p(x_{1}^{(k)}\mid x_{t})\leftarrow\frac{\exp(\log p(x_{1}^{(k)}\mid x_{t}))}{\sum_{k}\exp(\log p(x_{1}^{(k)}\mid x_{t}))}$ $\triangleright$ Normalize over $K$ components
29:end for
30:
31:Return Posterior $\{p(x_{1}\mid x_{t})\}$ for all time steps
Appendix E Power time schedule
Figure 13: Power distribution density
Figures 14 compares the time progression of $x_{t}$ between without and with power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ dataset. Figure 14(a) shows that without power time schedule, the samples remain nearly stationary until $t$ approaches $1$, while in Figure 14(b), with power time schedule, the samples start moving towards the target modes earlier, even at small $t$ values. This indicates that the power time schedule helps to alleviate the last-minute mode convergence issue discussed in Section 5.4 and Appendix D.
(a) Time progression of $x_{t}$ without power time schedule
(b) Time progression of $x_{t}$ with power time schedule
Figure 14: Test progression comparison between without and with power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ dataset.
E.1 Ablation study on skewness
We perform an ablation study on the skewness parameter of the power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ setting. Table 4 reports the Wasserstein-1 distance between the generated transformations and the ground truth group elements for different skewness values. Figure 15 visualizes the training curves for different skewness values. From the results, we observe that increasing the skewness of the power time schedule leads to better performance and faster convergence. Since the improvement saturates around skewness $5$, we use this value for all 3D experiments in the main paper.
Table 4: Ablation study on skewness of power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ setting. Wasserstein-1 distance between the generated transformations and the ground truth group elements are reported. Lower is better.
skewness 1 (uniform) 2 3 4 5
$0.328$ $0.171$ $0.137$ $0.121$ $0.118$
Figure 15: Ablation study on the skewness of power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ setting. $y$ axis denote the Wasserstein-1 distance between the generated transformations and the ground truth group elements, and $x$ axis denote training epoch.
Appendix F Additional Results
This section contains additional visualization results
F.1 2D arrow